# appAdditional References Appendix Ensuring Medical AI Safety: Explainable AI-Driven Detection and Mitigation of Spurious Model Behavior and Associated Data

### v1

@@ 3.3 Spatial Bias Localization
Notably, automated bias localization has the potential to reduce manual annotation efforts. These additional annotations can be utilized for various applications, e.g., as input for bias mitigation approaches requiring prior information on the bias to be unlearnt (see Sec. 5). Other applications include the design of metrics to measure artifact reliance or the spatial isolation of the bias to copy-paste it onto non-artifactual samples to measure the model’s sensitivity is towards the insertion of the artifact [59].
@@ 4 Concept Validation: Detecting Spurious Behavior
Given the large number of model parameters, detecting biased model representations can be like searching for a needle in a haystack, especially without prior knowledge of spurious correlations. To address this challenge, a common strategy is to identify outlier model behavior using a reference dataset. Automated detection approaches typically focus on either analyzing post-hoc explanations for a set of reference images to find anomalous model behavior [48, 27] or identifying outlier representations within the model itself [16, 53]. For concept validation (step ( ) in Fig. 1), we distinguish between the data perspective in Sec. 4.1, which focuses on detecting samples exhibiting outlier behavior, and the model perspective in Sec. 4.2, which aims to identify outlier concept representations within the model. However, it is to note that outlier model reasoning is not necessarily caused by spurious correlations, but can be (clinically) valid model behavior that is rarely used. Thus, detecting spurious correlations often requires manual inspection by human experts to determine whether outlier behavior is valid or caused by spurious correlations.
@@ 4.2 Model Perspective – Detecting Spurious Representations
In contrast to detecting spurious samples, recent work focuses on identifying spurious model internals directly. This aligns with mechanistic interpretability, which seeks to decipher the internal mechanics of DNNs [56, 31, 13]. As outlined in Sec. 3.1, various global XAIs methods aim to explain the role of individual neurons, and these insights can be leveraged to detect spurious model internals by clustering learned concepts and identifying outliers. Given input data with samples, Pahde et al. [59] compute spatially aggregated relevances . Subsequently, they compute the pairwise cosine distance per column (i.e., channel/concept) and embed the resulting distance matrix in a low-dimensional space using dimension reduction techniques like t-Distributed Stochastic Neighbor Embedding (t-SNE) [71] or Uniform Manifold Approximation and Projection (UMAP) [49]. This low-dimensional embedding can be visualized to identify outliers through human inspection or anomaly detection algorithms, such as the Local Outlier Factor [12]. In summary, outlier representations can be identified in an embedding representation , obtained as
where reduces the dimension to , and the pairwise distance function is applied along all channel dimensions in the latent representation , either given by activations or relevance scores for layer . Note that this approach assumes over-parameterization resulting in redundant neurons, allowing to distinguish between similar and dissimilar concept representations. An example is shown in Fig. 4 (bottom), where latent relevance scores from a ResNet50 model trained for melanoma detection are used to identify outlier concepts, specifically a cluster focusing on band-aids rather than clinically relevant features. Notably, Eq. 3 can easily be extended to find outlier directions instead of neurons. Specifically, this involves a linear transformation of latent representations using the directions of interest, e.g., obtained in unsupervised manner as described in Sec. 3.1.
Similarly, DORA embeds a pairwise distance matrix of neuron representations into 2D, but proposes a data-agnostic approach and a tailored distance function [16]. Specifically, they generate ActMax samples as concept representation for neurons, referred to as natural Activation-Maximization signals (n-AMS). Each neuron is represented by a representation activation vector , measuring how much other neurons activate on the given n-AMS, and compute pairwise distances between the vectors. The resulting distance matrix is embedded into lower dimension to identify outlier representations. Notably, instead of generating ActMax samples, this approach can also be applied on real samples from a reference dataset.
@@
5 Bias Mitigation
After detecting model biases and identifying biased samples, we aim to unlearn undesired behaviors through bias mitigation. A first line of approaches modifies the training data, e.g., by removing or manipulating biased samples, followed by retraining the model [77, 75]. While effectively mitigating biases, this method requires access to the full training set, can be costly, and may ignore valuable information, leading to practical limitations. In an iterative model development life cycle like Reveal2Revise, with continuous model validation and bias mitigation, full re-training is often infeasible. Thus, we focus on efficient bias mitigation approaches, that either finetune the (biased) model with a custom bias mitigation loss function or modify the model post-hoc without additional training. As such, Right for the Right Reason (RRR) [63] penalizes the alignment between the input gradient, i.e., the gradient of the prediction w.r.t. the input features, and ground truth masks localizing the artifact. Note, that this approach requires pixel-wise annotations for bias localization, which can be costly. However, (semi-)automated data annotation techniques outlined in Secs. 3.2 and 3.3 can reduce the manual effort.
As an alternative to input-level bias representations, the Class Artifact Compensation (ClArC) framework [3] models biases in latent space using CAVs, requiring only binary labels per sample to indicate the presence of the artifact. Inspired by RRR, Right Reason ClArC (RR-ClArC) [28] penalizes the feature use measured by the latent gradient pointing into the direction of the bias, as modeled via the CAV. Another research direction is training-free post-hoc model editing [3, 53, 8]. For instance, Projective ClArC (P-ClArC) [3] removes activations in the bias direction modeled via CAVs during the prediction. However, bias mitigation approaches risk “collateral damage”, meaning that whereas the biased concept is successfully suppressed, valid concepts entangled with the biased concept might be negatively impacted as well. To address this, reactive P-ClArC (rP-ClArC) [5] only targets samples meeting certain conditions, such as containing the artifact according to the CAV. A detailed description of bias mitigation approaches is provided in Appendix A.3.
@@ 6 Experiments
We evaluate our framework with four medical datasets from two modalities, namely vision and time-series. After describing the experimental setup (Sec. 6.1), we demonstrate the capabilities of our framework for bias identification (Sec. 6.2), the detection of biased samples (Sec. 6.3), bias localization (Sec. 6.4) and mitigation (Sec. 6.5).
@@ 6.1 Experimental Setup
The considered datasets include ISIC2019 for melanoma detection [19, 70, 20], HyperKvasir for the identification of gastrointestinal abnormalities [10], CheXpert with chest radiographs [46], and the PTB-XL dataset [73] with 12-lead ECG (time series) data. All vision datasets contain real-world artifacts that DNNs may utilize as spurious correlation, i.e., features unrelated to the task, yet correlating with the target label. ISIC2019 is particularly known for various artifacts like colorful band-aids near benign lesions and rulers or skin markers beside malignant lesions [62, 3, 18, 59]. Moreover, HyperKvasir contains insertion tubes predominantly in samples without abnormal conditions, while CheXpert samples with cardiomegaly contain pacemakers in radiographs more frequently than in healthy patients [75]. In addition, we insert controlled artifacts into a subset of images from exactly one class per dataset. Specifically, we insert a microscope-like artifact into melanoma samples in ISIC2019. Moreover, following Dreyer et al. [28], we insert a timestamp into disease-samples from HyperKvasir, mimicking timestamps added by scanning devices. For CheXpert, we increase the brightness of radiographs with cardiomegaly, while for PTB-XL, we insert a static noise into the first second of one lead for samples with left ventricular hypertrophy (LVH). Inserting these artifacts into of samples from exactly one class creates spurious correlations for that class. Further dataset details are provided in Appendix A.4 and examples of the artifacts are shown in Fig. 5.
@@ 6.2 Concept Validation: Identification of Spurious Behavior
Given the fitted models, we apply bias identification methods introduced in Sec. 4 to detect the expected spurious model behavior. Throughout this section, we focus on the ResNet50 model trained on ISIC2019 with various confounders. We further limit our analyses to samples from the melanoma class to prevent that clustering model behavior results in clusters representing different classes, and instead allowing us to identify spurious sub-strategies for predicting the considered class.
@@
We first apply SpRAy in input and latent space, computing input feature importance scores using LRP summed over color channels. To obtain latent relevances, we use intermediate relevance scores in the LRP computation process after the (out of four) residual block, max-pooling over spatial dimensions to yield relevance scores for layer with channels. The clustering of pairwise cosine distances between heatmaps is shown in Fig. 6 (top left). Detected outlier clusters contain samples with spatially coherent biases, e.g., the black circle around the lesions originating from microscopes (). In contrast, clustering latent relevance scores reveals more complex, less spatially dependent clusters, as shown in Fig. 6 (top right), including the ruler artifact () and the microscope (). Compared to those in input space, the cluster for the microscope in latent space represents a more diverse high-level concept. In the input space visualization, we further highlight samples from the ruler cluster detected in latent space. Instead of forming a cluster, they spread across the entire embedding space, indicating that the bias is too complex to be detected in input space.
@@
Next, we apply bias identification approaches from the model perspective by identifying outlier neurons based on activation pattern via DORA and relevance pattern by clustering pair-wise cosine distances between concept relevance scores. We focus on latent activations and relevances after the third residual block. DORA uses a distance function based on how neurons activate upon each others n-AMS, achieving high similarity when neurons activate upon similar input signal. A 2D visualization of the resulting distance matrix is shown in Fig. 6 (bottom left). Identified outlier concepts include ruler () and (white) hair (). We further compute pairwise cosine distances between latent relevance scores , aggregated over spatial dimensions, and apply UMAP to embed the resulting distance matrix in . This results in high similarity between neurons (concepts) that the model uses similarly for predictions. The concept clustering is visualized in Fig. 6 (bottom right), highlighting two outlier clusters focused on rulers () and blueish tint ().
Results for other classes, model architectures, and datasets are presented in Appendix A.6.1. This includes experiments with ECG data in Fig. 10, revealing the artificially inserted static noise in the attacked lead from both data and model perspectives. Notably, dominant spurious concepts, such as the artificial timestamp in HyperKvasir or the static noise in PTB-XL, may not be detected as outlier concepts. In such cases, analyzing prediction sub-strategies via PCX may provide additional insights on spurious inlier behavior. Hard-to-interpret concept representations pose another challenge for the model perspective. For example, the brightness artifact in CheXpert is not clearly visible in the concept UMAP (see. Fig. 15, right), but can easily be detected using SpRAy (Fig. 15, left) or PCX (Fig. 22). In summary, while all considered spurious features are detected, the choice of bias identification approach is crucial, as some shortcuts are easier to detect as outlier concept (e.g., ruler) and others via PCX (e.g., brightness, static noise in ECG).
@@ 6.4 Spatial Bias Localization
To spatially localize biases in input space with CAVs, we compute local explanations for the element-wise product of latent activations and concept direction (see Eq. 2). We use the controlled artifacts, specifically timestamp (HyperKvasir) and micropscope (ISIC2019) with ground truth concept localization masks for evaluation. We compute (1) the percentage of relevance within the ground truth mask and (2) the Jaccard index, or Intersection over Union (IoU), of the binarized localization mask with the ground truth. In Fig. 9, we report both metrics using CAVs computed on different layers of VGG16 and ResNet50. The layer choice for concept representations is key, as for example middle layers perform better to localize timestamps and earlier layers are more effective to localize the microscope. In comparison with Fig.7, we find that the optimal layer for bias localization may differ from the one for sample retrieval. Interestingly, the IoU for the microscope artifact is consistently low, as models primarily focus on the border of the circle instead of the entire area, as indicated by qualitative results in Appendix A.6.3. Unlike artifact relevance, the IoU metric also measures how much of the expected areas the computed mask does not cover.
@@ 6.5 Bias Mitigation
We unlearn detected biases using the methods summarized in Sec. 5. We utilize RRR as input-gradient-based bias mitigation approach for localizable artifacts and the ClArC framework for all artifacts, representing biases in latent space with CAVs. Specifically, we consider the penalty-based approach RR-ClArC and the training-free model editing methods P-ClArC and rP-ClArC. To measure the bias mitigation effect, we compute several metrics inspired by prior work [59]. First, we compare the accuracy on a clean (bias-free) and a biased test set, where the bias is artificially inserted into samples from all classes. Models impacted by spurious correlations are expected to perform poorly on the biased test set. In addition, we measure the model’s sensitivity towards the bias concept by computing (1) the percentage of relevance, measured via LRP, on the artifact region using ground truth masks, and (2) the TCAV score [47]. The latter is reported as , where 0 indicates no sensitivity and higher values reliance on the artifact. Low scores are preferred after bias mitigation. The results are compared to a Vanilla model that is finetuned without a bias mitigation loss term. In Tab. 1, we report results for ResNet50 models in the controlled settings with ISIC2019, HyperKvasir, and CheXpert. For RRR, we use ground truth bias localization masks and refrain from reporting results for CheXpert, as we consider the brightness artifact unlocalizable in input space. The results confirm that all models initially rely on the spurious correlation, indicated by a large gap between clean and biased accuracies for Vanilla models. All bias mitigation approaches improve the accuracy on the biased test set while maintaining high accuracy on the clean test set, demonstrating reduced reliance on the targeted biases. This is supported by decreased artifact relevance and . Qualitative results visualizing the decrease in artifact reliance with input relevance heatmaps and additional quantitative results for other model architectures and ECG data are provided in Appendix A.6.4.

### v2

@@ 3.3 Feature-Level Bias Annotation via Spatial Localization
Notably, automated bias localization has the potential to substantially reduce manual feature-level annotation efforts. These additional annotations can be utilized for various applications, e.g., as input for bias mitigation approaches requiring prior information on the bias to be unlearnt (see Sec. 5). Other applications include the design of metrics to measure artifact reliance or the spatial isolation of the bias to copy-paste it onto non-artifactual samples to measure the model’s sensitivity is towards the insertion of the artifact [57].
@@ 4 Concept Validation: Detecting Spurious Behavior
Given the large number of model parameters, detecting biased model representations can be like searching for a needle in a haystack, especially without prior knowledge of spurious correlations. To address this challenge, this section reviews and categorizes interpretability-driven approaches for detecting biased model behavior. A common strategy is to identify outlier model behavior using a reference dataset. Automated detection approaches typically focus on either analyzing post-hoc explanations for a set of reference images to find anomalous model behavior [48, 28] or identifying outlier representations within the model itself [17, 52]. For concept validation (step ( ) in Fig. 1), we distinguish between the data perspective in Sec. 4.1, which focuses on detecting samples exhibiting outlier behavior, and the model perspective in Sec. 4.2, which aims to identify outlier concept representations within the model. However, it is to note that outlier model reasoning is not necessarily caused by spurious correlations, but can be (clinically) valid model behavior that is rarely used. Thus, detecting spurious correlations often requires manual inspection by domain experts to determine whether outlier behavior is valid or caused by spurious correlations.
@@ 4.2 Model Perspective – Detecting Spurious Representations
In contrast to detecting spurious samples, recent work focuses on identifying spurious model internals directly. This aligns with mechanistic interpretability, which seeks to decipher the internal mechanics of DNNs [55, 32, 14]. As outlined in Sec. 3.1, various global XAI methods aim to explain the role of individual neurons, and these insights can be leveraged to detect spurious model internals by clustering learned concepts and identifying outliers. Given input data with samples, Pahde et al. [57] compute spatially aggregated relevances . Subsequently, they compute the pairwise cosine distance per column (i.e., channel/concept) and embed the resulting distance matrix in a low-dimensional space using dimension reduction techniques like t-Distributed Stochastic Neighbor Embedding (t-SNE) [72] or Uniform Manifold Approximation and Projection (UMAP) [49]. This low-dimensional embedding can be visualized to identify outliers through human inspection or anomaly detection algorithms, such as the Local Outlier Factor [13]. In summary, outlier representations can be identified in an embedding representation , obtained as
where reduces the dimension to with , and the pairwise distance function is applied along all channel dimensions in the latent representation , either given by activations or relevance scores for layer . Note that this approach assumes over-parameterization resulting in redundant neurons, allowing to distinguish between similar and dissimilar concept representations. An example is shown in Fig. 4 (bottom), where latent relevance scores from a ResNet50 model trained for melanoma detection are used to identify outlier concepts, specifically a cluster focusing on band-aids rather than clinically relevant features. Notably, Eq. 3 can easily be extended to find outlier directions instead of neurons. Specifically, this involves a linear transformation of latent representations using the directions of interest, e.g., obtained in unsupervised manner as described in Sec. 3.1.
Similarly, Data-agnOstic Representation Analysis (DORA) embeds a pairwise distance matrix of neuron representations into 2D, but proposes a data-agnostic approach and a tailored distance function [17]. This involves generating ActMax samples as concept representation for neurons, referred to as natural Activation-Maximization signals (n-AMS). Each neuron is represented by a representation activation vector , measuring how much other neurons activate on the given n-AMS, and compute pairwise distances between the vectors. The resulting distance matrix is embedded into lower dimension to identify outlier representations. Notably, instead of generating ActMax samples, this approach can also be applied on real samples from a reference dataset.
@@
5 Bias Mitigation Methods
After detecting and annotating biases, our goal is to unlearn undesired behaviors. Therefore, we briefly review existing bias mitigation approaches in this section (see Appendix A.4 for a more detailed overview). A first line of approaches modifies the training data, e.g., by removing or manipulating biased samples, followed by retraining the model [78, 76]. While effectively mitigating biases, this method requires access to the full training set, can be costly, and may ignore valuable information, leading to practical limitations. In an iterative model development life cycle like Reveal2Revise, with continuous model validation and bias mitigation, full re-training is often infeasible. Thus, we focus on efficient bias mitigation approaches, that either finetune the (biased) model with a custom loss function or modify the model post-hoc without additional training. As such, Right for the Right Reason (RRR) [63] penalizes the alignment between the input gradient, i.e., the gradient of the prediction w.r.t. the input features, and ground truth masks localizing the artifact.
As an alternative to input-level bias representations, the Class Artifact Compensation (ClArC) framework [3] models biases in latent space using CAVs, requiring only binary labels per sample to indicate the presence of the artifact. Inspired by RRR, Right Reason ClArC (RR-ClArC) [29] penalizes the feature use measured by the latent gradient pointing into the direction of the bias, as modeled via the CAV. Another research direction is training-free post-hoc model editing [3, 52, 9]. For instance, Projective ClArC (P-ClArC) [3] removes activations in the bias direction modeled via CAVs during the prediction. However, bias mitigation approaches risk “collateral damage”, meaning that whereas the biased concept is successfully suppressed, valid concepts entangled with the biased concept might be negatively impacted as well. To address this, reactive P-ClArC (rP-ClArC) [5] only targets samples meeting certain conditions, such as containing the artifact according to the CAV.
@@ 6 Experiments
We evaluate the extended Reveal2Revise framework with four medical datasets from two modalities, namely vision and time-series. We describe the experimental setup (Sec. 6.1) and demonstrate the capabilities of the framework for bias identification (Sec. 6.2), the detection of biased samples (Sec. 6.3), bias localization (Sec. 6.4) and mitigation (Sec. 6.5).
@@ 6.1 Experimental Setup
The considered datasets include ISIC2019 for melanoma detection [20, 71, 21], HyperKvasir for the identification of gastrointestinal abnormalities [11], CheXpert with chest radiographs [46], and the PTB-XL dataset [74] with 12-lead ECG (time series) data. All vision datasets contain real-world artifacts that DNNs may utilize as spurious correlation, i.e., features unrelated to the task, yet correlating with the target label. ISIC2019 is particularly known for various artifacts like colorful band-aids near benign lesions and rulers or skin markers beside malignant lesions [62, 19, 57]. Moreover, HyperKvasir contains insertion tubes predominantly in samples without abnormal conditions, while CheXpert samples with cardiomegaly contain pacemakers in radiographs more frequently than in healthy patients [76]. In addition, we insert controlled artifacts into a subset of images from exactly one class per dataset. Specifically, we insert a microscope-like artifact into melanoma samples in ISIC2019. Moreover, following Dreyer et al. [29], we insert a timestamp into disease-samples from HyperKvasir, mimicking timestamps added by scanning devices. For CheXpert, we increase the brightness of radiographs with cardiomegaly, while for PTB-XL, we insert a static noise into the first second of one lead for samples with left ventricular hypertrophy (LVH). Inserting these artifacts into of samples from exactly one class creates spurious correlations for that class. Further dataset details are provided in Appendix A.5 and examples of the artifacts are shown in Fig. 5.
@@ 6.2 Concept Validation: Identification of Spurious Behavior
Given fitted models, we apply bias identification methods described in Sec. 4 to detect the spurious model behavior. Throughout this section, we focus on the ResNet50 model trained on ISIC2019 with various confounders. We further limit our analyses to samples from the melanoma class to prevent that clustering model behavior results in clusters representing different classes, and instead allowing us to identify spurious sub-strategies for predicting the considered class. Results for other classes, model architectures, and datasets, including ECG data, are presented in Appendix A.7.1.
@@
We first apply SpRAy in input and latent space, computing input feature importance scores using LRP summed over color channels. To obtain latent relevances, we use intermediate relevance scores from the LRP computation after the (out of four) residual block, max-pooled over spatial dimensions to yield relevance scores for layer with channels. The clustering of pairwise cosine distances between heatmaps is shown in Fig. 6 (top left). Detected outlier clusters contain samples with spatially coherent biases, e.g., the black circle around the lesions originating from microscopes (). In contrast, clustering latent relevance scores reveals more complex, less spatially dependent clusters, as shown in Fig. 6 (top right), including the ruler artifact () and the microscope (). Compared to those in input space, the cluster for the microscope in latent space represents a more diverse high-level concept. In the input space visualization, we further highlight samples from the ruler cluster detected in latent space. Instead of forming a cluster, they spread across the entire embedding space, indicating that the bias is too complex to be detected in input space.
@@
Next, we apply bias identification approaches from the model perspective by identifying outlier neurons based on activation pattern via DORA and relevance pattern by clustering pair-wise cosine distances between concept relevance scores. We focus on latent activations and relevances after the third residual block. DORA uses a distance function based on how neurons activate upon each others n-AMS, achieving high similarity when neurons activate upon similar input signal. A 2D visualization of the resulting distance matrix is shown in Fig. 6 (bottom left). Identified outlier concepts include ruler () and (white) hair (). We further compute pairwise cosine distances between spatially aggregated latent relevance scores and apply UMAP to embed the resulting distance matrix in . This results in high similarity between neurons (concepts) that the model uses similarly for predictions. The concept clustering is visualized in Fig. 6 (bottom right), highlighting two outlier clusters focused on rulers () and blueish tint ().
The additional experiments with ECG data (see Appendix A.7.1) reveal the artificially inserted static noise in the attacked lead from both data and model perspectives. Moreover, additional vision experiments show that dominant spurious concepts, such as the artificial timestamp in HyperKvasir or the static noise in PTB-XL, may not be clearly detected as outlier concepts. In such cases, analyzing prediction sub-strategies via PCX may provide additional insights on spurious inlier behavior. Hard-to-interpret concept representations pose another challenge for the model perspective. For example, the brightness artifact in CheXpert is not clearly visible in the concept UMAP (see. Fig. 15, right), but can easily be detected using SpRAy (Fig. 15, left) or PCX (Fig. 22). In summary, while all considered spurious features are detected, the choice of bias identification approach is crucial, as some shortcuts are easier to detect as outlier concept (e.g., ruler) and others via PCX (e.g., brightness, static noise in ECG).
@@ 6.4 Spatial Bias Localization
To spatially localize biases in input space with CAVs, we compute local explanations for the element-wise product of latent activations and concept direction (see Eq. 2). We use the controlled artifacts, i.e., timestamp (HyperKvasir) and micropscope (ISIC2019) with ground truth concept localization masks for evaluation. We compute (1) the percentage of relevance within the ground truth mask and (2) the Jaccard index, or Intersection over Union (IoU), of the binarized mask with the ground truth. In Fig. 9, we report both metrics using CAVs computed on different layers of VGG16 and ResNet50. The layer choice for concept representations is key, as for example middle layers perform better to localize timestamps and earlier layers are more effective to localize the microscope. In comparison with Fig. 7, we find that the optimal layer for bias localization may differ from the one for sample retrieval. Interestingly, the IoU for the microscope artifact is consistently low, as models primarily focus on the border of the circle instead of the entire area, as indicated by qualitative results in Appendix A.7.3. Unlike artifact relevance, the IoU metric also measures how much of the expected areas the computed mask does not cover.
@@ 6.5 Bias Mitigation
We unlearn the detected biases using the methods summarized in Sec. 5. We utilize RRR as input-gradient-based bias mitigation approach for localizable artifacts and the ClArC framework for all artifacts, representing biases in latent space with CAVs. For the latter, we consider the penalty-based approach RR-ClArC and the training-free model editing methods P-ClArC and rP-ClArC. To measure the bias mitigation effect, we compute several metrics suggested by the Reveal2Revise framework. First, we compare the accuracy on a clean (bias-free) test set and a biased test set, where the bias is artificially inserted into samples from all classes. Models impacted by spurious correlations are expected to perform poorly on the biased test set. In addition, we measure the model’s sensitivity towards the bias concept by computing (1) the percentage of relevance, measured via LRP, on the artifact region using ground truth masks, and (2) the TCAV score [47]. The latter measures how sensitive the model is towards the artifact, represented as CAV, and is computed as the fraction of predictions that were positively influenced by the concept (see Appendix A.7.4 for details). It is reported as , where 0 indicates no sensitivity and higher values reliance on the artifact. Low scores are preferred after bias mitigation. The results are compared to a Vanilla model that is finetuned without a bias mitigation loss term. In Tab. 1, we report results for ResNet50 models in the controlled settings with ISIC2019, HyperKvasir, and CheXpert. For RRR, we use ground truth bias localization masks and refrain from reporting results for CheXpert, as we consider the brightness artifact unlocalizable in input space. The results confirm that all models initially rely on the spurious correlation, indicated by a large gap between clean and biased accuracies for Vanilla models. All bias mitigation approaches improve the accuracy on the biased test set while maintaining high accuracy on the clean test set, demonstrating reduced reliance on the targeted biases. This is supported by decreased artifact relevance and . Qualitative results visualizing the decrease in artifact reliance with input relevance heatmaps and additional quantitative results for other model architectures and ECG data are provided in Appendix A.7.4.